# GroundingSuite: Measuring Complex Multi-Granular Pixel Grounding

### v1

@@ 6.2 Comparisons with Other Automated Datasets
For fair comparisons, we randomly selected 100k samples from both GranD and the proposed GSTrain-10M to train models under the EVF-SAM framework while incorporating the RefCOCO series datasets. This experimental setting allows us to directly assess the quality and effectiveness of the two datasets while keeping other variables constant.
@@ Results.
As shown in Tab. 5, we evaluate models trained on each dataset across three standard benchmarks: RefCOCO, gRefCOCO, and RefCOCOm. The results reveal interesting performance patterns that highlight the relative strengths of each approach.
Data GSEval RefCOCO gRefCOCO RefCOCOm
GranD [21] 60.0 76.0 65.6 37.9
Ours 74.9
+14.9
76.7
+0.7
64.8
-0.8
39.5
+1.6
@@
On the standard RefCOCO benchmark, the proposed GSTrain-10M achieved a performance gain of 0.7 percentage points over GranD (76.7 vs. 76.0). More significantly, on the challenging RefCOCOm benchmark, which features more complex linguistic descriptions, our approach demonstrated a substantial improvement of 1.6 points (39.5 vs. 37.9).

### v2

@@ 6.2 Comparisons with Other Automated Datasets
For fair comparisons, we randomly selected 100k samples from both GranD and the proposed GSTrain-10M to train models under the EVF-SAM and LISA framework.
@@ Results.
Methods Dataset RefCOCO gRefCOCO RefCOCOm
EVF-SAM GranD 60.7 28.0 31.5
EVF-SAM GSTrain 63.5
+2.8
28.9
+0.9
32.7
+1.2
LISA-7B GranD 63.1 29.5 29.6
LISA-7B GSTrain 65.6
+2.5
31.3
+1.8
37.6
+8.0
@@
The result in Tab. 5 demonstrates that our GSTrain dataset consistently outperforms GranD across all benchmarks. With EVF-SAM, GSTrain improves performance by 2.8, 0.9, and 1.2 percentage points on RefCOCO, gRefCOCO, and RefCOCOm respectively. The gains are even more significant with LISA-7B, showing a 2.5 point improvement on RefCOCO and a remarkable 8.0 point gain on RefCOCOm (37.6 vs. 29.6), indicating GSTrain’s particular effectiveness with complex linguistic expressions.